# Weak-Strong Uniqueness and Extreme Wall Events at High Reynolds Number

### v1

@@ Abstract
Singular or weak solutions of the Euler equations have been hypothesized to acccount for anomalous
dissipation at very high Reynolds numbers and, in particular, to explain the d’Alembert
paradox of non-vanishing drag. A possible objection to this explanation is the mathematical
property called
“weak-strong uniqueness”, which requires that any admissable weak solution
of the Euler equations must coincide with the smooth Euler solution for the same initial data.
As an application of the Josephson-Anderson relation, we sketch a proof of conditional
weak-strong uniqueness for the potential Euler solution of d’Alembert within the class of strong inviscid limits.
We suggest that the mild conditions required for weak-strong uniqueness are, in fact, physically violated by violent
eruption of very thin boundary layers. We discuss observational signatures of these extreme events and
explain how the small length-scales involved could threaten the validity of a hydrodynamic description.
@@ II What is Weak-Strong Uniqueness?
One important implication of weak-uniqueness is on existence of blow-up or finite-time singularities for smooth Euler solutions. Indeed, weak-strong uniqueness for the inviscid limit solution implies that viscous energy dissipation must necessarily vanish over any finite time interval for which the smooth Euler solution exists Eyink (2008); Bardos and Titi (2013). Conversely, appearance of anomalous energy dissipation over a finite time interval for a flow with smooth initial data requires that the Euler solution with that initial data cannot remain smooth over the interval. A typical example of such initial data is the Taylor-Green vortex in periodic domain Taylor and Green (1937), although existing evidence for anomalous dissipation in this flow Fehn et al. (2022) seems less compelling to us than corresponding evidence for wall-bounded flows. In any case, a finite-time Euler singularity can be rigorously ruled out for certain smooth initial data, e.g. for any smooth data in a 2-dimensional spatial domain and for the stationary potential data in the 3-dimensional d’Alembert flow. Thus, finite-time blow-up of smooth Euler solutions cannot be the general route to turbulence and anomalous dissipation.
@@ III The New D’Alembert Paradox
However, a new d’Alembert paradox arises if one considers the limit There is substantial empirical evidence that drag remains non-vanishing in this limit, for example, the apparent non-zero asymptotic values of drag coefficients for solid bodies of various shapes. Nevertheless, any such limits must be described by generalized Euler solutions Lions (1996); DiPerna and Majda (1987) or, under reasonable assumptions, by weak Euler solutions as conjectured by Onsager Drivas and Nguyen (2019) and the admissability condition (II.1) furthermore must hold. Weak-strong uniqueness would require that the generalized Euler solution obtained in the limit must coincide with the smooth Euler solution of d’Alembert, suggesting that time-average drag must vanish in the limit To argue for this conclusion, one may consider the situation where the fluid starts at rest with and then smoothly accelerates to a velocity bounded in time. Taking so large that , then weak-strong uniqueness implies Quan and Eyink (2025a) that
@@ IV Weak-Strong Uniqueness Can Fail With Solid Walls
This is the famous condition shown by Kato Kato (1984) to be equivalent to strong convergence of the Navier-Stokes solution to a strong Euler solution, when the initial data also converge strongly and as long as the strong Euler solution exists. Since such strong limits of Navier-Stokes solutions are admissible weak Euler solutions, Kato’s theorem can be considered one of the earliest results of weak-strong uniqueness type. Another class of conditions sufficient for weak-strong uniqueness involve some continuity of velocity for the admissible weak Euler solution near the boundary. This was first proved by Bardos et al. (2014) who required that the velocity field be Hölder continuous in some neighborhood with and then proved by Wiedemann (2018) assuming only continuity in . Recently, we have shown Quan and Eyink (2025a) that even less stringent conditions on the weak Euler solution are sufficient for weak-strong uniqueness within the general class of admissible weak Euler solutions, namely:
@@
Note that where the “essential supremum” is the least upper bound of values such that the set has positive measure. The latter condition (IV.4) can be interpreted as uniform continuity at the boundary of the normal velocity component, since We refer to (IV.3)-(IV.4) as the “Drivas-Nguyen conditions” since they were first employed in Drivas and Nguyen (2018) to study anomalous energy dissipation in wall-bounded flow.
@@ V Proof Sketch for D’Alembert Flow
We thus use a version of a standard “relative energy” argument to prove weak-strong uniqueness Wiedemann (2018). In this context, our sufficient condition for weak-strong uniqueness is the vanishing only of the streamwise component of the skin friction, which is somewhat weaker than the condition invoked in Bardos and Titi (2013).
@@
Here is the anomalous energy dissipation and the last two terms are the inviscid limit of the Josephson-Anderson relation for power dissipated by drag, which appears as a source term in the rotational energy balance. Much of the technical work of the proof involves the justification of the energy balance equation (V.2), by smearing a corresponding local balance equation with a test function which is a smoothed version of the characteristic function restricted to and then taking the limit
@@
This argument shows that any strong- inviscid limit with initial data must in fact coincide with for a.e. times .
The applicability of this result to a smoothly accelerated body with is straightforward, assuming the validity of the incompressible Navier-Stokes model, since The relevance to the impulsively accelerated body requires more detailed justification because of the singular nature of that problem. It is worthwhile to discuss here since it raises some physics issues which will be important later.
@@
Some doubts remain, however, whether our analysis applies to a physical realization of impulsive acceleration in a laboratory experiment. The kinetic description of a gas by the Boltzmann equation in the low-density limit is now believed to be incomplete, missing stochastic effects of molecular fluctuations Spohn (1981); Bouchet (2020). Such fluctuation effects do not necessarily vitiate Onsager’s theoretical description of high Reynolds number flows by weak Euler solutions Eyink and Peng (2024). However, to our knowledge, weak-strong uniqueness results have not yet been proved for inviscid limits of models incorporating fluctuations. Thus, it seems advisable to focus theoretical treatment and future empirical investigation on the more regular (and commonplace) problem of gradually and smoothly accelerated bodies. Rather surprisingly, however, we shall find that even the problem of smooth acceleration may encounter such difficulties, because breakdown of weak-strong uniqueness requires such extreme events that the validity of a macroscopic hydrodynamic description again is threatened. See Section VII.3.
@@ VI.1 First Alternative:
Under this condition, weak-strong uniqueness holds. As a consequence, for initial data strongly converging to pure potential, drag on a body vanishes over any fixed time-interval as see (III.4). If the drag vanishes sufficiently slowly, however, then it might be very difficult to distinguish empirically from drag strictly non-vanishing. The phenomenon of very slowly vanishing drag and dissipation has been termed a “weak dissipative anomaly” Bedrossian et al. (2019). Such weak anomalies seem to occur physically in a number of wall-bounded flows, such as circular pipe flows with perfectly smooth walls (see Eyink (2024a) for a summary). Strong anomalous dissipation and drag may furthermore occur experimentally if the initial data is not pure potential in the high-Reynolds number limit. Although the rotational flow from a thin viscous boundary layer vanishes in the energy norm as there may in addition be small but non-vanishing vorticity in the incoming flow. This is similar to what occurs in so-called “bypass transition”, where small levels of turbulent fluctuations in the background flow can trigger laminar-to-turbulent transition. In order to discuss both of the above possibilities more quantitatively, we can apply the analysis of the previous section.
@@ VII Extreme Wall Events
We have already discussed in Section IV three rather modest-looking conditions that suffice for weak-strong uniqueness. Within the class of weak-* limits of Navier-Stokes solutions with bounded energy in general domains, Bardos and Titi (2013) have shown that the condition (IV.1) of vanishing skin-friction suffices for weak-strong uniqueness of any smooth Euler solution. Furthermore, Bardos and Titi (2013) have shown that the condition (IV.2) of vanishing energy dissipation in a “Kato layer” implies the previous condition (IV.1). In fact, the third set of conditions (IV.3),(IV.4) on continuity of normal velocity at the wall implies also the vanishing skin-friction condition (IV.1) for strong inviscid limits of Navier-Stokes solutions. The proof in Quan and Eyink (2025b) is based on the analogy between skin friction and viscous energy dissipation, on the one hand, and energy cascade through scales and momentum cascade through space, on the other hand Jiménez (2012). Thus, non-vanishing skin friction can be understood as a “strong momentum anomaly” and the continuity conditions (IV.3),(IV.4) are analogous to the Onsager condition on velocity Hölder exponent which forbids an energy to small scales. The analogous idea of the proof in Quan and Eyink (2025b) is that non-vanishing velocity toward the wall, at any positive distance, is required to carry momentum to the viscous sink at the wall in the infinite Reynolds number limit. Thus, a strong momentum anomaly is possible only if the normal velocity is discontinuous at the wall (or, more accurately, not uniformly continuous).
Somewhat surprisingly, the conditions for a momentum anomaly and breakdown of weak-strong uniqueness are much more severe than those needed for energy cascade, requiring essentially a shock-like discontinuity at the wall, or an Hölder singularity! Interestingly, this possibility seems to have been anticipated by G. I. Taylor as early as 1915, who wrote the following in discussing interactions of atmospheric turbulence with the solid ground:
@@
The Kato condition (IV.2) likewise involves an extreme situation of non-vanishing viscous dissipation in a shock-like layer at a wall. Kato-type boundary layers of thickness it is worth noting, were proposed already in 1923 by Burgers as a mechanism to produce anomalous drag Burgers (1923)111Burgers’ discussion was longer than Taylor’s Taylor (1915) but his arguments, as he admitted himself, were “in many respects vague”. He proceeded by deriving integral mean balances for momentum and energy and by investigating the parameter values to maximize drag. He correctly concluded that a boundary layer with thickness would produce skin friction non-vanishing as Unfortunately, this argument leads to non-vanishing drag coefficient for smooth walls, whereas Burgers had noted himself that only “for rough-walled tubes is approximately independent” of .. The corresponding rigorous conditions for weak-strong uniqueness, although originally derived for internal flows in bounded domains, apply also to external flows around bodies: see Section V.
@@ VII.1.1 Skin Friction
The probability distribution of wall shear stress has been a focus for previous numerical studies Lenaers et al. (2012); Srinivas et al. (2024), in particular the streamwise component which contributes to drag. Note that it is also non-vanishing of the streamwise component which is relevant for possible physical violation of weak-strong uniqueness for plug-flow in a channel, with In fact, since all of the theorems about infinite- limit cited in this paper require non-dimensionalization of flow variables in outer units, what is relevant is scaled by Since the results in Lenaers et al. (2012); Srinivas et al. (2024) are instead scaled by the friction-velocity squared, we have the scanned the data in Figure 2 of Lenaers et al. (2012) and rescaled the results by the factor The parameterization of the Prandtl-Kármán drag law from equation (12) of Zanoun et al. (2021) was used to determine from and hence the ratio The results for the probability density function of in outer units are plotted in Figure 1, at three values of the Reynolds number
@@ VII.1.3 Dissipation in a Viscous Near-Wall Layer
The Kato condition (IV.2) for channel flow is just the requirement that the net energy dissipation integrated over the viscous sublayer and buffer layer should vanish relative to as Although there have been a great many studies of viscous dissipation in turbulent channel flows, we are not aware of any study of this particular issue. The prediction of the Prandtl-Kármán theory is that dissipation in the buffer layer and viscous sublayer should scale as smaller than the total dissipation in the log-layer by a factor of about Both, however, tend to zero relative to as These considerations suggest that the Kato condition should be satisfied and indeed convergence to plug flow has been argued to occur in smooth-walled, plane-parallel channel flows Cantwell (2019).
@@ VII.2 Smooth Vortices Impinging on a Wall
Most importantly, the paper Nguyen van yen et al. (2018) presented strong evidence also for the extreme events that are required in order to violate weak-strong uniqueness. Figure 12a of Nguyen van yen et al. (2018) plotted versus Reynolds number the maximum vorticity at the wall, which was found to scale as before the blow-up of the Prandtl solution but as a short time after the blow up. Since skin friction is related to wall vorticity by for stick b.c., the above observation corresponds to as at least pointwise. Furthermore, Figure 12b of Nguyen van yen et al. (2018) plotted also versus Reynolds number the enstrophy which was likewise found to scale as before the Prandtl blow-up but as afterward. Because total viscous energy dissipation can be written as the latter scaling implies a dissipative anomaly. No evidence was provided in Nguyen van yen et al. (2018) for violation of the Drivas-Nguyen conditions (IV.3), (IV.4), but the authors did note “… a blow-up of the wall-normal velocity associated with an abrupt acceleration of fluid particles away from the wall” (p.697). Although further confirmation would be desirable, the paper Nguyen van yen et al. (2018) presents to our knowledge the most complete evidence for physical violation of weak-strong uniqueness within the class of inviscid limits.

### v3

@@ Abstract
Singular or weak solutions of the incompressible Euler equations have been hypothesized to account for anomalous dissipation at very high Reynolds numbers and, in particular, to explain the d’Alembert paradox of non-vanishing drag. A possible objection to this explanation is the mathematical property called “weak-strong uniqueness”, which requires that any admissable weak solution of the Euler equations must coincide with the smooth Euler solution for the same initial data. As an application of the Josephson-Anderson relation, we sketch a proof of conditional weak-strong uniqueness for the potential Euler solution of d’Alembert within the class of strong inviscid limits. We suggest that the mild conditions required for weak-strong uniqueness are, in fact, physically violated by violent eruption of very thin boundary layers. We discuss observational signatures of these extreme events and explain how the small length-scales involved could threaten the validity of a hydrodynamic description.
@@ II What is Weak-Strong Uniqueness?
One important implication of weak-strong uniqueness is on existence of blow-up or finite-time singularities for smooth Euler solutions. Indeed, weak-strong uniqueness for the inviscid limit solution implies that viscous energy dissipation must necessarily vanish over any finite time interval for which the smooth Euler solution exists Eyink (2008); Bardos and Titi (2013). Conversely, appearance of anomalous energy dissipation over a finite time interval for a flow with smooth initial data requires that the incompressible Euler solution with that initial data cannot remain smooth over the interval. A typical example of such initial data is the Taylor-Green vortex in periodic domain Taylor and Green (1937), although existing evidence for anomalous dissipation in this flow Fehn et al. (2022) seems less compelling to us than corresponding evidence for wall-bounded flows. In any case, a finite-time Euler singularity can be rigorously ruled out for certain smooth initial data, e.g. for any smooth data in a 2-dimensional spatial domain and for the stationary potential data in the 3-dimensional d’Alembert flow. Thus, finite-time blow-up of smooth Euler solutions cannot be the general route to turbulence and anomalous dissipation in incompressible fluid flows.
@@ III The New D’Alembert Paradox
However, a new d’Alembert paradox arises if one considers the limit Stewartson (1981). There is substantial empirical evidence that drag remains non-vanishing in this limit, for example, the apparent non-zero asymptotic values of drag coefficients for solid bodies of various shapes. Nevertheless, any such limits must be described by generalized Euler solutions Lions (1996); DiPerna and Majda (1987) or, under reasonable assumptions, by weak Euler solutions as conjectured by Onsager Drivas and Nguyen (2019) and a suitable generalization of the admissability condition (II.1) furthermore must hold. Weak-strong uniqueness would require that the generalized Euler solution obtained in the limit must coincide with the smooth Euler solution of d’Alembert, suggesting that time-average drag must vanish in the limit To argue for this conclusion, one may consider the situation where the fluid starts at rest with and then smoothly accelerates to a velocity bounded in time. Taking so large that , then weak-strong uniqueness implies Quan and Eyink (2025a) that
@@ IV Weak-Strong Uniqueness Can Fail With Solid Walls
This is the famous condition shown by Kato Kato (1984) to be equivalent to strong convergence of the Navier-Stokes solution to a strong Euler solution, when the initial data also converge strongly and as long as the strong Euler solution exists. Since such strong limits of Navier-Stokes solutions are admissible weak Euler solutions, Kato’s theorem can be considered one of the earliest results of weak-strong uniqueness type. A more recent work of Kelliher Kelliher (2017) has comprehensively discussed these various conditions for inviscid limits of closed flows in spatially bounded domains and shown, in particular, that the condition (IV.1) can be weakened to
where is an assumed strong Euler solution. Another class of conditions sufficient for weak-strong uniqueness involve some continuity of velocity for the admissible weak Euler solution near the boundary. This was first proved by Bardos et al. (2014) who required that the velocity field be Hölder continuous in some neighborhood with and then proved by Wiedemann (2018) assuming only continuity in . Recently, we have shown Quan and Eyink (2025a) that even less stringent conditions on the weak Euler solution are sufficient for weak-strong uniqueness within the general class of admissible weak Euler solutions, namely:
@@
Note that where the “essential supremum” is the least upper bound of values such that the set has positive measure. The latter condition (IV.5) can be interpreted as uniform continuity at the boundary of the normal velocity component, since We refer to (IV.4)-(IV.5) as the “Drivas-Nguyen conditions” since they were first employed in Drivas and Nguyen (2018) to study anomalous energy dissipation in wall-bounded flow.
@@ V Proof Sketch for D’Alembert Flow
We thus use a version of a standard “relative energy” argument to prove weak-strong uniqueness Wiedemann (2018). In this context, our sufficient condition for weak-strong uniqueness is that for all times, which is nearly the same condition (IV.3) invoked by Kelliher for closed flows in bounded domains.
@@
Here is the anomalous energy dissipation and the last two terms are the inviscid limit of the Josephson-Anderson relation for power dissipated by drag, which appears as a source term in the rotational energy balance. Much of the technical work of the proof involves the justification of the energy balance equation (V.2), by smearing a corresponding local balance equation with a test function which is a smoothed version of the characteristic function where is the ball of radius centered at the origin, restricted to and then taking the limit
@@
This argument shows that any strong- inviscid limit with initial data must in fact coincide with for a.e. times . The applicability of this result to a smoothly accelerated body with is straightforward, assuming the validity of the incompressible Navier-Stokes model, since The relevance to the impulsively accelerated body requires more detailed justification because of the singular nature of that problem. It is worthwhile to discuss here since it raises some physics issues which will be important later.
@@
Some doubts remain, however, whether our analysis applies to a physical realization of impulsive acceleration in a laboratory experiment. The kinetic description of a gas by the Boltzmann equation in the low-density limit is now believed to be incomplete, missing stochastic effects of molecular fluctuations Spohn (1981); Bouchet (2020); Sawant et al. (2021). Such fluctuation effects do not necessarily vitiate Onsager’s theoretical description of high Reynolds number flows by weak Euler solutions Eyink and Peng (2024). However, to our knowledge, weak-strong uniqueness results have not yet been proved for inviscid limits of models incorporating fluctuations. Thus, it seems advisable to focus theoretical treatment and future empirical investigation on the more regular (and commonplace) problem of gradually and smoothly accelerated bodies. Rather surprisingly, however, we shall find that even the problem of smooth acceleration may encounter such difficulties, because breakdown of weak-strong uniqueness requires such extreme events that the validity of a macroscopic hydrodynamic description again is threatened. See Section VII.3.
@@ VI.1 First Alternative:
Under this condition, weak-strong uniqueness holds. As a consequence, for initial data strongly converging to pure potential, drag on a body vanishes over any fixed time-interval as see (III.4). If the drag vanishes sufficiently slowly, however, then it might be very difficult to distinguish empirically from drag strictly non-vanishing. The phenomenon of very slowly vanishing drag and dissipation has been termed a “weak dissipative anomaly” Bedrossian et al. (2019). Such weak anomalies seem to occur physically in a number of wall-bounded flows, such as circular pipe flows with perfectly smooth walls (see Eyink (2024a) for a summary). Convergence to a smooth potential Euler solution has been suggested to occur also in some bluff body flows, e.g. rapidly rotating cylinders where separation is argued to be suppressed Glauert (1957), although recent high Reynolds number simulations do not seem to support that conclusion Aljure et al. (2015). Strong anomalous dissipation and drag may furthermore occur experimentally if the initial data is not pure potential in the high-Reynolds number limit. Although the rotational flow from a thin viscous boundary layer vanishes in the energy norm as there may in addition be small but non-vanishing vorticity in the incoming flow. This is similar to what occurs in so-called “bypass transition”, where small levels of turbulent fluctuations in the background flow can trigger laminar-to-turbulent transition. In order to discuss both of the above possibilities more quantitatively, we can apply the analysis of the previous section.
@@ VII Extreme Wall Events
We have already discussed in Section IV three rather modest-looking conditions that suffice for weak-strong uniqueness. Within the class of weak-* limits of Navier-Stokes solutions with bounded energy in general domains, Bardos and Titi (2013) have shown that the condition (IV.1) of vanishing skin-friction suffices for weak-strong uniqueness of any smooth Euler solution. Furthermore, Bardos and Titi (2013) have shown that the condition (IV.3) of vanishing energy dissipation in a “Kato layer” implies the previous condition (IV.1). In fact, the third set of conditions (IV.4),(IV.5) on continuity of normal velocity at the wall implies also the vanishing skin-friction condition (IV.1) for strong inviscid limits of Navier-Stokes solutions. The proof in Quan and Eyink (2025b) is based on the analogy between skin friction and viscous energy dissipation, on the one hand, and energy cascade through scales and momentum cascade through space, on the other hand Jiménez (2012). Thus, non-vanishing skin friction can be understood as a “strong momentum anomaly” and the continuity conditions (IV.4),(IV.5) are analogous to the Onsager condition on velocity Hölder exponent which forbids an energy to small scales. The analogous idea of the proof in Quan and Eyink (2025b) is that non-vanishing velocity toward the wall, at any positive distance, is required to carry momentum to the viscous sink at the wall in the infinite Reynolds number limit. Thus, a strong momentum anomaly is possible only if the normal velocity is is not uniformly continuous at the wall.
Somewhat surprisingly, the conditions for a momentum anomaly and breakdown of weak-strong uniqueness are much more severe than those needed for energy cascade, requiring essentially a shock-like “discontinuity” at the wall, or an Hölder singularity! Interestingly, this possibility seems to have been anticipated by G. I. Taylor as early as 1915, who wrote the following in discussing interactions of atmospheric turbulence with the solid ground:
@@
The Kato condition (IV.3) likewise involves an extreme situation of non-vanishing viscous dissipation in a shock-like layer at a wall. Kato-type boundary layers of thickness it is worth noting, were proposed already in 1923 by Burgers as a mechanism to produce anomalous drag Burgers (1923)111Burgers’ discussion was longer than Taylor’s Taylor (1915) but his arguments, as he admitted himself, were “in many respects vague”. He proceeded by deriving integral mean balances for momentum and energy and by investigating the parameter values to maximize drag. He correctly concluded that a boundary layer with thickness would produce skin friction non-vanishing as Unfortunately, this argument leads to non-vanishing drag coefficient for smooth walls, whereas Burgers had noted himself that only “for rough-walled tubes is approximately independent” of .. The corresponding rigorous conditions for weak-strong uniqueness, although originally derived for internal flows in bounded domains, apply also to external flows around bodies: see Section V.
@@ VII.1.1 Skin Friction
The probability distribution of wall shear stress has been a focus for previous numerical studies Lenaers et al. (2012); Srinivas et al. (2024), in particular the streamwise component which contributes to drag. Note that it is also non-vanishing of the streamwise component which is relevant for possible physical violation of weak-strong uniqueness for plug-flow in a channel, with a spacetime constant velocity in the streamwise or -direction. In fact, since all of the theorems about infinite- limit cited in this paper require non-dimensionalization of flow variables in outer units, what is relevant is scaled by Since the results in Lenaers et al. (2012); Srinivas et al. (2024) are instead scaled by the friction-velocity squared, we have scanned the data in Figure 2 of Lenaers et al. (2012) and rescaled the results by the factor The parameterization of the Prandtl-Kármán drag law from equation (12) of Zanoun et al. (2021) was used to determine from and hence the ratio The results for the probability density function of in outer units are plotted in Figure 1, at three values of the Reynolds number
@@ VII.1.3 Dissipation in a Viscous Near-Wall Layer
The Kato condition (IV.3) for channel flow is just the requirement that the net energy dissipation integrated over the viscous sublayer and buffer layer should vanish relative to as Although there have been a great many studies of viscous dissipation in turbulent channel flows, we are not aware of any study of this particular issue. The prediction of the Prandtl-Kármán theory is that dissipation in the buffer layer and viscous sublayer should scale as smaller than the total dissipation in the log-layer by a factor of about Both, however, tend to zero relative to as These considerations suggest that the Kato condition should be satisfied and indeed convergence to plug flow has been argued to occur in smooth-walled, plane-parallel channel flows Cantwell (2019).
@@ VII.2 Smooth Vortices Impinging on a Wall
Most importantly, the paper Nguyen van yen et al. (2018) presented strong evidence also for the extreme events that are required in order to violate weak-strong uniqueness. Figure 12a of Nguyen van yen et al. (2018) plotted versus Reynolds number the maximum vorticity at the wall, which was found to scale as before the blow-up of the Prandtl solution but as a short time after the blow up. Since skin friction is related to wall vorticity by for stick b.c., the above observation corresponds to as at least pointwise. Furthermore, Figure 12b of Nguyen van yen et al. (2018) plotted also versus Reynolds number the enstrophy which was likewise found to scale as before the Prandtl blow-up but as afterward. Because total viscous energy dissipation can be written as the latter scaling implies a dissipative anomaly. No evidence was provided in Nguyen van yen et al. (2018) for violation of the Drivas-Nguyen conditions (IV.4), (IV.5), but the authors did note “… a blow-up of the wall-normal velocity associated with an abrupt acceleration of fluid particles away from the wall” (p.697). Although further confirmation would be desirable, the paper Nguyen van yen et al. (2018) presents to our knowledge the most complete evidence for physical violation of weak-strong uniqueness within the class of inviscid limits.